Fairfax Paint operates stores in Virginia. The firm is evaluating the Vienna project, which would involve opening a new store in Vienna. During year 1, Fairfax Paint would have total revenue of 373,000 dollars and total costs of 292,000 dollars if it pursues the Vienna project, and the firm would have total revenue of 305,000 dollars and total costs of 271,000 if it does not pursue the Vienna project. Depreciation taken by the firm would be 73,000 dollars if the firm pursues the project and 40,000 dollars if the firm does not pursue the project. The tax rate is 15 percent. What is the relevant operating cash flow (OCF) for year 1 of the Vienna project that Fairfax Paint should use in its NPV analysis of the Vienna project?
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

A B C D E 1 Incremental With Vienna Project $373,000.00 $292,000.00 $73,000.00 Without Vienna project $305,000.00 $271,000.00

Cell reference -

With Vienna Project Without Vienna project Incremental =C3-D3 373000 292000 73000 305000 271000 40000 =C4-D4 =C5-D5 Revenue V

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.